Solid Ca(IO3)2 is added to 0.0100 M KIO3. Titration shows that the concentration of iodate in the solution is now 0.0150 M IO3-. a) What portion of the iodate concentration was derived from the dissolving of calcium iodate? b) What is the concentration of calcium created by the dissolving solid?
Answer #1

a) Portion of Iodate concentration derived from Ca(IO3)2 = Concentration of IO3- - Concentration of KIO3

=> 0.0150 - 0.0100

=> 0.0050M

b) So, on dissolving Ca(IO3)2, the concentration of IO3- is increased by 0.050M

On dissolving, Ca(IO3)2 will split as below

Ca(IO3)2 ================= Ca(2+) + 2IO3-

Concentration of calcium created = 1/2 * concentration increase of IO3- = 1/2 * 0.050 = 0.025 M